OpenStudy (anonymous):

∫ (e^(cos3x)) times (sin3xdx)

OpenStudy (dape):

The derivative of cos3x is -sin3x. So we try the substitution: \[ u=cos3x \\ du=-sin3x \ dx \] This yields: \[ \int{e^{cos3x}sin3x \ dx} = \int{e^u \ -du} = -e^u+C \] Substituting back for x we have the solution \[ -e^{cos3x}+C \]

OpenStudy (dape):

That is e^u times (-du) in the middle step.

OpenStudy (dape):

Ah, I'm tired. du is -3sin3x dx and not -sin3x. The solution is therefore \[ -\frac{1}{3}e^{cos3x}+C \]

OpenStudy (dape):

Sorry for that, just ask if my answer became confusing.

OpenStudy (anonymous):

Nope, explanation was great! thanks a ton